Ashley has twice as much money invested at 5% simple annual interest as she does at 4%. If her yearly income from the two investments is $182, ho wmuch does she have at each rate?
Answer by mananth(16949)

Ashley has twice as much money invested at 5% simple annual interest as she does at 4%. If her yearly income from the two investments is $182, ho wmuch does she have at each rate?
4%------x
5%------2x
..
Interest = 182
4x+5*2x=182*100
14x=18200
/14
x=$ 1300 @4 %
$ 2600 @5%